### v1

# bywithappendix

Abstract.
We identify all smooth manifolds of curves for Heath-Jarrow-Morton models that are consistent with any tangential diffusion coefficient. In fact, we show that these manifolds cannot be affine but must be of linear-rational type.
Keywords: Heath-Jarrow-Morton models, term structure models, invariant manifolds, finite-dimensional realisations
1. Introduction
In fixed-income markets, the basic financial instruments are securities that are based on interest rate payments. A common example among these are government bonds, which typically entail the holder to a fixed coupon, which is a regular payment, and an additional nominal at the end of the contract. The two most principle approaches to model all of the government bonds at once are the Heath-Jarrow-Morton (HJM) approach [HJM92] and the LIBOR market model [BGM97]. In the HJM approach an underlying infinite-dimensional process $(f(t,T))_{0\leq t\leq T}$ is modelled, where $f(t,T)$ is called the forward rate at time $t$ and time of maturity $T$. In this framework, the zero-coupon bond price at time $t$ with contract maturity date $T$ is defined via
$P(t,T)=e^{-\int_{t}^{T}f(t,s)ds}.$
It is convenient to use (with slight abuse of notation) the parametrisation $f_{t}(x):=f(t,t+x)$ due to [Mus93] which expresses the forward rate in terms of time and time-to-maturity instead of time and maturity. One can consider the forward rate as a curve-valued object and arrive at the Heath-Jarrow-Morton-Musiela (HJMM) formulation. The original HJM-drift condition which ensures absence of arbitrage has been reformulated in [Mus93] into the time-to-maturity formulation.
In order to make these models computable, one possibility is to set up a finite factor model $(Z_{t})_{t\geq 0}$ with $f_{t}=g(Z_{t},\cdot)$ for some deterministic function $g$. A more principled approach to this is discussed in [Bjö04] where, instead of a finite-parameter set of curves, a manifold $\mathcal{M}$ of curves is used. [Bjö04, p. 17] asks three questions which are all about the relation of a forward rate curve model $(f_{t})_{t\geq 0}$ and some manifold of curves $\mathcal{M}$, as for instance: Under which condition does a given forward rate curve model stay on a prescribed manifold of curves?
From a more general standpoint, manifolds which remain invariant under (stochastic) dynamical systems are an important object of study as a solution space of (stochastic) partial differential equations ((S)PDEs). In mathematical finance those invariant manifolds arise naturally when considering finite-dimensional realisations of (potentially infinite-dimensional) consistent term structure models (see, e.g. [BC99, BL00, BLS02, BS01, FT01, Tap05, Tap15, Tap10]). There are several motivations for considering finite-dimensional realisations. For instance, finite dimensional models are easier simulated in a computer and if there is a finitely parametrised family of curves it is easy to statistically decide for a curve from data than in some infinite dimensional case.
We are interested in the question of when, for a given manifold of curves $\mathcal{M}$, a model for the forward rate curve can be constructed. This problem was analysed in [Fil00], where equivalent conditions have been derived. In particular, any diffusion model for $f$ which is compatible with a given manifold $\mathcal{M}$ must have at least a diffusion coefficient which is tangential to $\mathcal{M}$. Additional results, as in e.g. [FTT14] further characterise invariant finite-dimensional manifolds for SPDEs describing the evolution of the forward curve $f$. Examples of finite-dimensional manifolds which are invariant under consistent HJMM dynamics are known. Most notably, affine term structures which are finite dimensional affine spaces have been studied in e.g. [FT03, FT04] where the authors show that the only HJMM-type models which admit a finite-dimensional realisation for essentially any initial curve $f_{0}$ are affine in the direction of the diffusion coefficient. In contrast to our problem specification, the authors construct invariant (finite-dimensional) manifolds given a diffusion coefficient, which turn out to be necessarily affine.
Under the assumption of a finitely parametrised set of curves $\mathcal{M}=\{g(z,\cdot):z\in\mathbb{R}^{d}\}$ with a smooth $g$, any $\mathbb{R}^{d}$-valued diffusion $Z$ satisfies $f_{t}:=g(Z_{t},\cdot)\in\mathcal{M}$ by construction and naturally has the property that the diffusion coefficient of $f$ is tangential to $\mathcal{M}$. In fact, the diffusion coefficient of $f$ can be easily represented via the diffusion coefficient of $dZ_{t}=\beta_{t}dt+\sigma_{t}dW_{t}$ through an application of Itô’s formula:
$df_{t}=(\partial_{1}g(Z_{t},\cdot)\beta_{t}+\frac{1}{2}\partial_{11}g(Z_{t},\cdot)\sigma_{t}^{2})dt+\partial_{1}g(Z_{t},\cdot)\sigma_{t}dW_{t}.$
Given a fixed a curve fitting method for identifying the underlying forward rate curve given market data, the possible set of curves $\mathcal{M}=\{g(z,\cdot):z\in\mathbb{R}^{d}\}$ is typically finitely parametrised. In a second step, one might want to estimate the implicit diffusion coefficient of the parameter process from data and use the result of the estimation for a model for $Z$. This would require that the diffusion coefficient obtained by the method in question is indeed compatible with the given manifold $\mathcal{M}$. Given these considerations, our main question is this:
  • (Q)
    Which smooth manifolds $\mathcal{M}$ of curves have consistent HJM-type models for any possible tangential diffusion coefficient?
The precise mathematical statement is summarised in Theorem 3.3 below.
The problem of calibrating financial models is well known in mathematical finance. In particular, calibration of term structure models has to be done in a careful way to preserve no-arbitrage conditions specified by the respective model. Recalibrating models frequently may lead to violations thereof. In [RT17] the authors consider discrete-time recalibration schemes which generate arbitrage-free models for any time via so-called forward characteristics. Applications for term structure models of bonds and volatility curves are considered. To calibrate the term structure of interest rates, often times parametric families of curves are calibrated to properly interpolate observed prices. It is well-known (see, e.g. [Fil01]) that these families have to be chosen carefully, as otherwise dynamic consistency fails to be maintained. More recent approaches have been studied, as in e.g. [WJ24], where the authors attempt to employ a smoothing procedure for cubic splines to fit yield curves. To maintain dynamic consistency, the authors add higher-degree polynomials to the linear span of functions. In [LMS24], the authors use deep-learning methods to obtain a low-dimensional representation of the term structures via autoencoders. These, however, do not necessarily stay on the invariant manifold generated by the consistent model and instead remains within a vicinity dependent on the volatility.
In this article, we characterise all invariant finite-dimensional manifolds for the HJMM-equation which are compatible with any possible tangential diffusion coefficient. Our main result is that such a manifold must have a special rational structure. As a corollary, we find that an affine term structure where all tangential diffusion coefficients are compatible consists of a single point and implies, in fact, a deterministic model.
The article is structured as follows: Section 2 provides a brief introduction to our setting in the context of invariant manifolds for interest rates. In Section 3 we formulate our main result. The proofs for our main statements are relegated to Section 4 which concludes the treatment of the topic. Appendix A contains technical tools we make use of throughout the paper.
2. Preliminaries
2.1. Notation
Throughout this paper, let $(\Omega,\mathfrak{A},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ be a complete and right-continuous filtered probability space supporting a $d$-dimensional standard Brownian motion $W$. Let $X$ and $Y$ be sets. We will denote by $\mathfrak{F}(X,Y)$ the space of functions from $X$ to $Y$. Furthermore, we will use the notation $\mathfrak{M}(X,Y)$ for the space of measurable functions from $X$ to $Y$, $\text{Lip}(X,Y)$ the space of Lipschitz-continuous functions from $X$ to $Y$, $C^{k}(X,Y)$ the space of $k$-times continuously (Fréchet) differentiable functions form $X$ to $Y$ and $L(X,Y)$ the space of linear functions from $X$ to $Y$. Let $(\mathcal{H},|\cdot|)$ be a separable Hilbert space of functions from $\mathbb{R}_{+}$ to $\mathbb{R}$, that is $\mathcal{H}\subseteq\mathfrak{F}(\mathbb{R}_{+},\mathbb{R})$ with
  • (H1)
    $\delta_{0}:\mathcal{H}\rightarrow\mathbb{R},h\mapsto h(0)$ is continuous linear.
  • (H2)
    $\mathcal{\operatorname{S}}_{h}:\mathcal{H}\rightarrow\mathcal{H},f\mapsto(x\mapsto f(x+h))$ defines a $c_{0}$-semigroup $\mathcal{(}\operatorname{S}_{h})_{h\geq 0}$ on $\mathcal{H}$ whose generator will be denoted by $\partial_{x}$, cf. [EK09, pp. 6–8].
We denote the scalar product on $\mathcal{H}$ by $\langle h_{1},h_{2}\rangle_{\mathcal{H}}$ for $h_{1},h_{2}\in\mathcal{H}$. We will make use of a Hilbert space $\mathcal{H}$ satisfying Assumptions LABEL:H1 and LABEL:H2 throughout the paper without explicitly referencing them. An example of a Hilbert space with those properties is the forward curve space $\mathcal{H}_{w}$ (see [Fil01]) defined as follows:
Definition 2.1.
Let $w:\mathbb{R}_{+}\rightarrow[1,\infty)$ be a non-decreasing $C^{1}$-function such that
$w^{-1/3}\in L^{1}(\mathbb{R}_{+}).$
Define the norm
$\lVert h\rVert_{w}^{2}:=\lvert h(0)\rvert^{2}+\int_{0}^{\infty}\lvert h^{\prime}(x)\rvert^{2}w(x)dx$
and set
[…]
We will denote the standard unit basis vectors in $\mathbb{R}^{d}$ as $e_{i}$ for $i=1,\dots,d$ and the $d$-dimensional identity matrix as $\mathbbm{1}_{d}$. The standard scalar product on $\mathbb{R}^{d}$ will be denoted by $\langle v,w\rangle$ for $v,w\in\mathbb{R}^{d}$. We will denote elements $v$ of $\mathbb{R}^{d}$ as a vector coordinate-wise in the form $v=(v_{1},\dots,v_{d})$. Additionally, for $v\in\mathbb{R}^{d_{1}}$ and $w\in\mathbb{R}^{d_{2}}$ we will write the vector $u\in\mathbb{R}^{d_{1}+d_{2}}$ obtained by concatenating $v$ with $w$ as $u=(v,w)$. Let $X$ and $Y$ be two (topological) sets, such that $0\in Y$. Given a function $f:X\rightarrow Y$, we will denote by $\operatorname{supp}(f)$ the support of the function $f$, that is the smallest closed set such that $f(x)=0$ for $x\in X\setminus\operatorname{supp}(f)$.
2.2. Setting
We are interested in characterising the finite dimensional manifolds $\mathcal{M}\subseteq\mathcal{H}$ which are invariant under the HJMM-equation relative to any diffusion coefficient which is tangential to the manifold. The motivation of the problem comes from a statistical consideration. If one has decided for a set of curves $\mathcal{M}$ and estimates the diffusion coefficient of a process running on $\mathcal{M}$, then we like to be sure to obtain a model which is free of arbitrage in the sense of NAFLVR (see [CKT16]). Indeed, it is desirable for a calibrated model to remain within the admissible space of functions no matter the values of the estimated parameters. We proceed by making mathematically clear what we mean and summarise our main finding in Theorem 3.3 below. We will assume that the manifold is embedded into a Hilbert space of functions which is also used to formalise the HJMM-equation.
Assume now we are given a stochastic basis $(\Omega,\mathfrak{A},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ and a Hilbert space $\mathcal{H}$ fulfilling Assumptions LABEL:H1 and LABEL:H2. We begin with an observation.
Remark 2.2.
Note that the point evaluation $\delta_{x}:=\delta_{0}\operatorname{S}_{x}:\mathcal{H}\rightarrow\mathbb{R}$ is continuous linear for any $x\geq 0.$ Let $f\in\operatorname{dom}(\partial_{x})$. Then we have
$\partial_{x}f(x)$ $=\left(\lim_{h\searrow 0}\frac{\operatorname{S}_{h}f-f}{h}\right)(x)$
$=\lim_{h\searrow 0}\frac{f(x+h)-f(x)}{h}$
where the second equality follows from the continuity of $\delta_{x}$. This means that $\partial_{x}$ is the right-derivative and its domain contains only functions that have a right derivative and where the right-derivative defines an element of $\mathcal{H}$.
For the sake of completeness and to be mathematically precise, we provide a (classical) definition of manifolds we will rely on for the rest of the paper.
Definition 2.3.
Let $n\in\mathbb{N}$. A non-empty subset $\mathcal{M}\subseteq\mathcal{H}$ is called a $d$-dimensional manifold of class $C^{k}$ if for all $m\in\mathcal{M}$ there exists an open neighbourhood $V\subset\mathcal{H}$ of $m$, an open set $U\in\mathbb{R}^{d}$ and a map $\chi\in C^{k}(U,\mathcal{H})$ such that
  1. i)
    $\chi:U\rightarrow V\cap\mathcal{M}$ is a homeomorphism.
  2. ii)
    $D\chi(y)$ has rank $d$ for all $y\in U$.
The map $\chi$ is called a local parametrisation of $\mathcal{M}$. We shall denote by $T_{m}$ the tangent space of $\mathcal{M}$ at $m\in\mathcal{M}$, that is $T_{m}:=\{D\chi(y)v:v\in\mathbb{R}^{d}\}$ for $y\in U$ so that $\chi(y)=m$. For the exact definition we refer to [FTT14a, Section 3].
Since we are interested in invariant manifolds for consistent interest rate models, we now formally introduce the notion of the HJMM-equation. This will serve as a basis for further considerations in the theory. Indeed, we will tacitly assume that the stochastic processes we work with in the further parts of the paper satisfy the HJMM-equation.
Definition 2.4.
A stochastic process $f$ with values in $\mathcal{H}$ is a solution to the HJMM-equation with diffusion operator $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ if
  1. i)
    $\Sigma$ is measurable and locally bounded in the sense of [Tap12, Definition 3.5].
  2. ii)
    $f_{t}=\mathcal{\operatorname{S}}_{t}f_{0}+\int_{0}^{t}\operatorname{S}_{t-s}\beta(f_{s})ds+\int_{0}^{t}\operatorname{S}_{t-s}\Sigma(f_{s})dW_{s}$ for all $t\geq 0$, ,
where $\beta(h):=\frac{1}{2}\Sigma(h)\int_{0}^{\cdot}(\delta_{x}\Sigma(h))^{*}dx$ and we assume $\beta(h)\in\mathcal{H}$. We will say that $f$ has starting value $f_{0}$.
Remark 2.5.
If $f$ is a solution to the HJMM-equation with diffusion operator $\Sigma$, then it is a mild solution to
$df_{t}=(\partial_{x}f_{t}+\beta(f_{t}))dt+\Sigma(f_{t})dW_{t}$ (1)
in the sense of [DZ92, Section 6.1], where $\beta$ is as in Definition 2.4.
Following the arguments in [Zab00], the study of invariant manifolds is interesting for the sake of model calibration. Indeed, for curve fitting for interest rate models, it is desirable for the stochastic models to stay within a prescribed manifold while still staying consistent with the NAFLVR condition. This is made precise in the following definitions.
Definition 2.6.
[…]
Definition 2.7.
Let $\mathcal{M}\subset\mathcal{H}$ and define the parameter set $\Lambda\subset\mathfrak{M}(\mathcal{H},\mathcal{H})\times\mathfrak{M}(\mathcal{H},L(\mathbb{R}^{d},\mathcal{H}))$. We say $\mathcal{M}$ is fully invariant for $\Lambda$ if $\mathcal{M}$ is invariant under the SPDE
$df_{t}=(\partial_{x}f_{t}+\beta(f_{t}))dt+\Sigma(f_{t})dW_{t}$ (2)
for any pair $(\beta,\Sigma)\in\Lambda$.
We are mainly interested in solutions to the HJMM-equation, hence we will distinguish that special case. Furthermore, to ensure existence of solutions and sufficient regularity, we will impose some additional technical constraints on the parameter set. These are laid out in the following definition:
Definition 2.8.
Let $\mathcal{M}\subset\mathcal{H}$ be a submanifold and consider the HJMM-equation in Definition 2.4. Define $\mathcal{B}\subset C^{1}(\mathcal{H},L(\mathbb{R}^{d},\mathcal{H}))\cap\emph{Lip}(\mathcal{H},L(\mathbb{R}^{d},\mathcal{H}))$ to be a subset of functions $\Sigma$ with $\operatorname{supp}(\Sigma)\cap\mathcal{M}$ compact.
Define now the parameter set for the HJMM-equation as
$\Lambda_{\text{HJMM}}=\bigg\{(\beta,\Sigma)\;\bigg|\;\beta(h)=\frac{1}{2}\Sigma(h)\int_{0}^{\cdot}(\delta_{x}\Sigma(h))^{*}dx,\Sigma\in\mathcal{B}\bigg\}.$
We say that $\mathcal{M}$ is fully invariant under the HJMM-equation if it is fully invariant for $\Lambda_{\text{HJMM}}$.
Remark 2.9.
Let $\mathcal{M}$ be a $C^{2}$-manifold, $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ be of class $C^{1}$ and $\beta:\mathcal{H}\rightarrow\mathcal{H}$ locally Lipschitz and locally bounded in the sense of [Fil01, Definition 2.4.2, Definition 2.4.3], respectively. [Fil00, Theorem 3], [FT03] have shown that the following statements are equivalent:
  1. i)
    $\mathcal{M}$ is locally invariant under Equation 1 with drift $\beta$ and diffusion $\Sigma$.
  2. ii)
    The following statements hold:
    1. (a)
      $\mathcal{M}\subseteq\operatorname{dom}(\partial_{x})$.
    2. (b)
      $\Sigma(m)\in L(\mathbb{R}^{d},T_{m})$ for any $m\in\mathcal{M}$.
    3. (c)
      $\partial_{x}m+\beta(m)-\frac{1}{2}\sum_{j}D(\Sigma^{j})(m)\Sigma^{j}(m)\in T_{m}$ for any $m\in\mathcal{M}$, where $\Sigma^{j}(m):=\Sigma(m)e_{j}$ for $e_{j}\in\mathbb{R}^{d}$ the $j$-th unit basis vector of $\mathbb{R}^{d}$.
Moreover, [Fil00, Theorem 2] yields that if $\mathcal{M}$ is closed in $\mathcal{H}$, then invariance is ensured for $\tau=\infty$. [FTT14] extend this to a setting with jump diffusions with the additional requirement of $\mathcal{M}$ being of class $C^{3}$.
Remark 2.10.
A natural construction of a manifold $\mathcal{M}$ is as follows: Choose $g\in C^{2}(\mathbb{R}^{d},\mathcal{H})$ injective with $\operatorname{rank}(Dg(y))=d$ for any $y\in\mathbb{R}^{d}$ and define the $C^{2}$-manifold
$\mathcal{M}:=\{g(y)\;|\;y\in\mathbb{R}^{d}\}.$
If $f$ is a path-continuous solution to the HJMM-equation with diffusion coefficient some $\Sigma$, then it can be written as $f_{t}=g(Y_{t})$ for some path-continuous $\mathbb{R}^{d}$-valued diffusion process $Y$. For practical applications, it appears more convenient to start out with a process $Y$ with $dY_{t}=b_{t}dt+a_{t}dW_{t}$ and then define define $f_{t}:=g(Y_{t})$. Note that in this case, one obtains from Itô’s formula
$df_{t}=(Dg(Y_{t})b_{t}+\frac{1}{2}D^{2}g(Y_{t})(a_{t},a_{t}))dt+Dg(Y_{t})a_{t}dW_{t}$
and we see immediately that $\Sigma_{t}:=Dg(Y_{t})\in L(\mathbb{R}^{d},T_{f_{t}})$ for any $t\geq 0$. If we can rewrite $\Sigma_{t}=\Sigma(f_{t})$ for some measurable $\Sigma$, then this reads as $\Sigma(m)\in T_{m}$ for any $m\in\mathcal{M}$ (at least those $m$ which $f$ can attain). That is, the tangential condition 2.9.ii)b naturally holds from this perspective. The manifold $\mathcal{M}:=\{g(y):y\in\mathbb{R}^{d}\}$ is fully invariant under the HJMM-equation if and only if for any possible bounded choice of diffusion coefficient $a$ with compact support for $Y$ and any starting value $Y_{0}$ one can find a corresponding drift coefficient $b$ such that with $dY_{t}=b_{t}dt+a_{t}dW_{t}$ one has $g(Y)$ is a solution to the HJMM-equation.
3. Main results
In this section, we will formally state our main result, which gives a description of finite-dimensional manifolds which are fully invariant under the HJMM-equation in the sense of Definition 2.8.
Definition 3.1.
Let $\mathcal{M}$ be a $d$-dimensional $C^{2}$ manifold. We call $\mathcal{M}$ linear-rational if there is an open and connected set $U\subseteq\mathbb{R}^{d}$, and $C^{2}$-functions $c:\mathbb{R}_{+}\rightarrow\mathbb{R}$, $u:\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$ such that
$\mathcal{M}=\Bigg\{\frac{c^{\prime}+\sum_{j=1}^{d}z_{j}u^{\prime}_{j}}{1-\left(c+\sum_{j=1}^{d}z_{j}u_{j}\right)}\;\Bigg|\;z\in U\Bigg\}$
and where the denominators appearing in the equation for $\mathcal{M}$ are positive, i.e. we have
$c(x)+\sum_{i=1}^{d}z_{j}u_{j}(x)<1\qquad\text{for any }x\in\mathbb{R}_{+},z\in U.$
Example 3.2.
We now give an example of a linear-rational manifold $\mathcal{M}$ being a $C^{2}$-submanifold of a Hilbert space $\mathcal{H}$ satisfying Assumptions LABEL:H1 and LABEL:H2. Let $\alpha>0$ and choose $\lambda_{1}<\dots<\lambda_{d}<-\alpha/2$. Set $U:=(-\infty,0)^{d}$ and define the functions $u_{j}(x):=e^{\lambda_{j}x}$ for $j=1,\dots,d$ and $c\equiv 0$ and let $\mathcal{M}$ be given as in Definition 3.1. For any $f\in\mathcal{M}$, we have for $z_{1},\dots,z_{d}\in U$
$f(x)=\frac{\sum_{j=1}^{d}z_{j}\lambda_{j}e^{\lambda_{j}x}}{1-\sum_{j=1}^{d}z_{j}e^{\lambda_{j}x}}.$
and $f\in\mathcal{H}_{w}$ for $w(x):=e^{\alpha x}$, that is, $\mathcal{M}$ is a $C^{2}$-submanifold of $\mathcal{H}_{w}$, where $\mathcal{H}_{w}$ is the forward curve space defined in Definition 2.1.
Theorem 3.3.
Let $\mathcal{M}$ be a connected $d$-dimensional $C^{2}$-manifold such that $\mathcal{M}$ is fully invariant under the HJMM-equation. Then $\mathcal{M}$ is linear-rational in the sense of Definition 3.1
The proof of Theorem 3.3 is provided at the end of Section 4. Our main result implies that affine term structures are usually not fully invariant under the HJMM-equation. This is summarised in the next statement.
Corollary 3.4.
Let $\mathcal{M}\subseteq\mathcal{H}$ be a finite dimensional affine space and assume that it is fully invariant under the HJMM-equation. Then $\mathcal{M}$ contains a single point.
Proof.
Let $h_{0},h_{1}\in\mathcal{M}$ and define the function $f:\mathbb{R}_{+}\rightarrow\mathcal{H}$, $t\mapsto(1-t)h_{0}+th_{1}$. Since $\mathcal{M}$ is an affine space, $f$ is $\mathcal{M}$-valued. It then follows from Lemma A.5 that $\mathcal{M}$ is a singleton. ∎
Remark 3.5.
The difference between our setting and the setting in [FT03, FT04] is as follows. In [FT04, Theorem 5.3] they answer the question: Given an admissible volatility structure, which HJMM-models permit a finite-dimensional realisation for essentially any initial curve $f_{0}$. Indeed, any invariant, finite-dimensional manifold constructed for such an initial curve $f_{0}$ turns out to be affine in the direction of the volatility. Our result answers a different question, namely: Which manifolds $\mathcal{M}$ are invariant for any tangential diffusion coefficient $\Sigma$ and any initial curve $f_{0}\in\mathcal{M}$? Here, the construction already assumes that for any $f_{0}\in\mathcal{M}$, there is a solution for the HJMM-equation with diffusion $\Sigma$ which is $\mathcal{M}$-valued, and our result reveals the structure of these manifolds. It turns out that these are of a linear-rational nature, cf. Theorem 3.3.
The next theorem characterises the curves which appear in Theorem 3.3 in the description of the manifold $\mathcal{M}$.
Theorem 3.6.
Let $\mathcal{H}^{0}\subseteq\mathcal{H}$ be a subspace such that for any $h\in\mathcal{H}^{0}$ we have that the function $f_{h}(x):=h(x)\int_{0}^{x}h(u)du\in\mathcal{H}$ for $x\geq 0$.
Let $\mathcal{M}\subseteq\mathcal{H}^{0}$ be a connected $d$-dimensional $C^{2}$-submanifold which is closed in $\mathcal{H}$. Then the following statements are equivalent:
  1. i)
    $\mathcal{M}$ is fully invariant under the HJMM-equation.
  2. ii)
    $\mathcal{M}$ is linear-rational in the sense of Definition 3.1 and there is a matrix $M\in\mathbb{R}^{(d+1)\times(d+1)}$ such that the functions $c$ and $u$ fulfill $(c,u)(x)=(\mathbbm{1}_{d+1}-e^{xM})e_{1}$. In particular, $c$ and $u$ are real analytic functions.
The proof of Theorem 3.6 is given at the end of Section 4.
Remark 3.7.
The Hilbert space $\mathcal{H}_{w}$ given in Definition 2.1 with $\mathcal{H}_{w}^{0}\subseteq\mathcal{H}_{w}$ given by [Fil01, Equation 5.12] satisfies that $f_{h}$ as given in the previous theorem is in $\mathcal{H}_{w}$ for any $h\in\mathcal{H}_{w}^{0}$ as shown in [Fil01, Lemma 5.2.1].
4. Proof of the main statement
Throughout this section, let $\operatorname{J}$ denote the operator defined in Lemma A.2. We consider the transformation
$\Psi:$ $\mathcal{H}\rightarrow C(\mathbb{R}_{+},\mathbb{R}),$ (3)
$f\mapsto 1-\exp(-\operatorname{J}_{\cdot}f).$
Lemma 4.1.
The map $\Psi$ from Equation 3 has the following properties:
  1. i)
    $\Psi f\in C^{1}(\mathbb{R}_{+},\mathbb{R})$ for any $f\in\mathcal{H}$.
  2. ii)
    $(\delta_{x}\circ\Psi)\in C^{\infty}(\mathcal{H},\mathbb{R})$ for any $x\geq 0$.
  3. iii)
    The map $\tilde{\Psi}:\mathcal{H}\times\mathbb{R}_{+}\rightarrow\mathbb{R},(f,x)\mapsto(\Psi f)(x)$ satisfies $\tilde{\Psi}\in C^{\infty,1}(\mathcal{H}\times\mathbb{R}_{+},\mathbb{R})$.
  4. iv)
    $(\Psi^{-1}h)(x)=\frac{h^{\prime}(x)}{1-h(x)}$ for any $h\in\operatorname{ran}(\Psi)$ and any $x\geq 0$.
Proof.
Let $\Psi$ be given as in Equation 3.
  1. i)
    Let $h\in\mathcal{H}$ be arbitrary. By the properties of the exponential map and the integral, we have
    $(\Psi f)^{\prime}(x)=f(x)\exp(-\operatorname{J}_{x}f),$
    which is a continuous map.
  2. ii)
    The integral $\operatorname{J}_{x}$ is a linear operator on $\mathcal{H}$ for any $x\geq 0$, and the exponential map is smooth on $\mathcal{H}$, whence the statement follows.
  3. iii)
    This follows directly.
  4. iv)
    Let $h\in\operatorname{ran}(\Psi)$. Then there is $f\in\mathcal{H}$ with
    $h(x)=1-\exp(-\operatorname{J}_{x}f),\quad x\geq 0,$
    and hence
    $\frac{h^{\prime}(x)}{1-h(x)}=f(x),\quad x\geq 0$
    as required.
If $f_{t}$ is the forward rate curve seen at a single point of time $t$, then the corresponding zero-coupon bond prices are given by $P(t,t+x)=\exp(-\int_{0}^{x}f_{t}(s)ds)$. The bond prices are also given via the discount curve $h$ by $P(t,t+x)=1-h_{t}(x)$ and we see that $h_{t}(x)=1-P(t,t+x)=1-\exp(-\int_{0}^{x}f_{t}(s)ds)=(\Psi f_{t})(x)$. Furthemore, the short rate process $r_{t}:=\lim_{T\rightarrow T}-\partial_{T}\log(P(t,T))$ fulfills $r_{t}=\delta_{0}(f_{t})$. The next lemma characterises the transformation of the dynamics of $f$ into the dynamics of the transformed process $\Psi f$ in dependence of its state.
Lemma 4.2.
Let $\Sigma:\mathcal{H}\rightarrow L(\mathcal{H},\mathbb{R}^{d})$ be locally bounded and measurable and let $f$ be an $\mathcal{H}$-valued progressively measurable process satisfying
$f_{t}=\operatorname{S}_{t}f_{0}+\int_{0}^{t}\operatorname{S}_{t-s}\beta(f_{s})ds+\int_{0}^{t}\operatorname{S}_{t-s}\Sigma(f_{s})dW_{s}\quad\text{for any }t\geq 0,$
where $\beta(h):=\frac{1}{2}\Sigma(h)\operatorname{J}_{\cdot}\Sigma(h)^{*}$. Define for any $t\geq 0$ the process
$h_{t}:=\Psi f_{t}.$
Then we have
$h_{t}=S_{t}h_{0}+\int_{0}^{t}\operatorname{S}_{t-s}\beta^{h}(h_{s})ds+\int_{0}^{t}\operatorname{S}_{t-s}\Sigma^{h}(h_{s})dW_{s},$ (4)
where
$\beta^{h}(h)$ $=(h-1)\delta_{0}(\Psi^{-1}h),$ (5)
$\Sigma^{h}(h)$ $=(1-h)\operatorname{J}_{\cdot}\Sigma(\Psi^{-1}h)$
Moreover, one has
$\Sigma(f)(x)=\partial_{x}\left(\exp(\operatorname{J}_{x}f)\Sigma^{h}(\Psi f)\right)$ (6)
Proof.
Let $r_{t}=\delta_{0}(f_{t})$ for any $t\geq 0$ and define for $0\leq t\leq T$
$\tilde{h}_{t}(T-t):=(\Psi f_{t})(T-t)=1-\exp\left(-\operatorname{J}_{T-t}f_{t}\right).$
We have
$\operatorname{J}_{T-t}(f_{t})$ $=(\operatorname{J}_{T}-\operatorname{J}_{t})f_{0}+\int_{0}^{t}(\operatorname{J}_{T-s}-\operatorname{J}_{t-s})\beta(f_{s})+\int_{0}^{t}(\operatorname{J}_{T-s}-\operatorname{J}_{t-s})\Sigma(f_{s})dW_{s}$ (7)
$=\operatorname{J}_{T}f_{0}+\int_{0}^{t}(\operatorname{J}_{T-s})\beta(f_{s})+\int_{0}^{t}(\operatorname{J}_{T-s})\Sigma(f_{s})dW_{s}$
$-\left(\operatorname{J}_{t}f_{0}+\int_{0}^{t}\operatorname{J}_{t-s}\beta(f_{s})ds+\int_{0}^{t}\operatorname{J}_{t-s}\Sigma(f_{s})dW_{s}\right).$
Furthermore, we find that the short rate satisfies
$\int_{0}^{t}r_{x}dx$ $=\int_{0}^{t}\left(\delta_{0}\operatorname{S}_{x}f_{0}+\int_{0}^{x}\delta_{0}\operatorname{S}_{x-s}\beta(f_{s})ds+\int_{0}^{x}\delta_{0}\operatorname{S}_{x-s}\Sigma(f_{s})dW_{s}\right)dx$
$=\int_{0}^{t}\delta_{x}f_{0}dx+\int_{0}^{t}\int_{0}^{x}\delta_{x-s}\beta(f_{s})dsdx+\int_{0}^{t}\int_{0}^{x}\delta_{x-s}\Sigma(f_{s})dW_{s}dx$
Using Fubini’s theorem to exchange the order of integration we find that
$\int_{0}^{t}r_{x}dx$ $=\int_{0}^{t}f_{0}(x)dx+\int_{0}^{t}\int_{0}^{t-s}\delta_{x}\beta(f_{s})dxds+\int_{0}^{t}\int_{0}^{t-s}\delta_{x}\Sigma(f_{s})dxdW_{s}$ (8)
$=\operatorname{J}_{t}f_{0}+\int_{0}^{t}\operatorname{J}_{t-s}\beta(f_{s})ds+\int_{0}^{t}\operatorname{J}_{t-s}\Sigma(f_{s})dW_{s}$
[…]
$\operatorname{J}_{T-t}(f_{t})=\operatorname{J}_{T}f_{0}+\int_{0}^{t}\operatorname{J}_{T-s}\beta(f_{s})+\int_{0}^{t}\operatorname{J}_{T-s}\Sigma(f_{s})dW_{s}-\int_{0}^{t}r_{s}ds.$
Consequently, we have by Itô’s formula
$\tilde{h}_{t}(T-t)=$ $\tilde{h}_{0}(T-t)$
$+\int_{0}^{t}\exp(-\operatorname{J}_{T-s}f_{s})\left(\operatorname{J}_{T-s}\beta(f_{s})-r_{s}-\frac{1}{2}\operatorname{J}_{T-s}(\Sigma(f_{s})\Sigma(f_{s})^{*}\right)ds$
$+\int_{0}^{t}\exp(-\operatorname{J}_{T-s}f_{s})\operatorname{J}_{T-s}\Sigma(f_{s})dW_{s}.$
After substituting the expression for the HJM-drift for $\beta$, we observe after integrating that $\operatorname{J}_{\cdot}\beta(f)=\frac{1}{2}\operatorname{J}_{\cdot}\left(\Sigma(f)\Sigma(f)^{*}\right)$. Thus, we find
$\tilde{h}_{t}(T-t)=\tilde{h}_{0}(T-t)-\int_{0}^{t}\exp(-\operatorname{J}_{T-s}f_{s})r_{s}ds+\int_{0}^{t}\exp(\operatorname{J}_{T-s}f_{s})\operatorname{J}_{T-s}\Sigma(f_{s})dW_{s}.$
Defining the drift- and diffusion coefficient as in Equation 5, we have $\tilde{h}_{t}(x)=\delta_{x}h_{t}$ for any $t\geq 0$ and $x\geq 0$, where $h$ satisfies Equation 4 and the claim follows. Finally, Equation 6 follows by inverting the identity for $\Sigma^{h}$ in Equation 5. ∎
To proceed with our proof, we need an additional result for existence of solutions to the HJMM-equation in the special case of the regularity assumptions we have outlined in Definition 2.8. As it turns out, this can be deduced from standard results.
Lemma 4.3.
Let $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ be Lipschitz, such that $\emph{supp}(\Sigma)\subseteq B$ for some bounded set $B\subset\mathcal{H}$. Define $\beta(h):=\frac{1}{2}\Sigma(h)\int_{0}^{\cdot}(\delta_{x}\Sigma(h))^{*}dx$. Then there is an $\mathcal{H}$-valued stochastic process $(f_{t})_{t\geq 0}$ such that
$f_{t}=\operatorname{S}_{t}f_{0}+\int_{0}^{t}\operatorname{S}_{t-s}\beta(f_{s})ds+\int_{0}^{t}\operatorname{S}_{t-s}\Sigma(f_{s})dW_{s},$
for any $t\geq 0$.
Proof.
First, observe that for any $h\in\mathcal{H}$
$\begin{aligned} \beta(h)&=\frac{1}{2}\Sigma(h)\int_{0}^{\cdot}(\delta_{x}\Sigma(h))^{*}dx=\frac{1}{2}\Sigma(h)\int_{0}^{\cdot}\Sigma(h)^{*}\delta_{x}^{*}dx\\ &=\frac{1}{2}\Sigma(h)\Sigma(h)^{*}\int_{0}^{\cdot}\delta_{x}^{*}dx.\end{aligned}.$
Since $\mathcal{H}$ fulfills Assumption LABEL:H1, it is an RKHS and therefore there exists for all $x\in\mathbb{R}_{+}$ a function $k_{x}\in\mathcal{H}$ such that for $h\in\mathcal{H}$ we have $\delta_{x}h=\langle k_{x},h\rangle_{\mathcal{H}}$. We may now identify $\delta_{x}^{*}$ by using the reproducing property. We have for $h\in\mathcal{H}$ and $y\in\mathbb{R}$
$\langle h,\delta_{x}^{*}y\rangle_{\mathcal{H}}=\langle\delta_{x}h,y\rangle=y\langle h,k_{x}\rangle_{\mathcal{H}}=\langle h,yk_{x}\rangle_{\mathcal{H}}.$
Thus, $\delta_{x}^{*}$ can be identified with the function $y\mapsto yk_{x}$. We have $k_{x}=1\cdot k_{x}\delta_{x}^{*}(1)=\operatorname{S}^{*}_{x}\delta_{0}^{*}(1)=\operatorname{S}^{*}_{x}(1\cdot k_{0})=\operatorname{S}^{*}_{x}k_{0}$. Consider for $h,h^{\prime}\in\mathcal{H}$ the map $\phi_{h,h^{\prime}}:\mathbb{R}_{+}\rightarrow\mathbb{R}$
$x\mapsto\langle h,\operatorname{S}^{*}_{x}h^{\prime}\rangle_{\mathcal{H}}=\langle\operatorname{S}_{x}h,h^{\prime}\rangle_{\mathcal{H}}.$ (9)
Since $\mathcal{H}$ fulfills Assumption LABEL:H2, $\operatorname{S}$ is strongly continuous. This implies that the maps $\phi_{h,h^{\prime}}$ are continuous for all $h,h^{\prime}i\in\mathcal{H}$. This implies that the adjoint semigroup $(\operatorname{S}_{x}^{*})_{x\geq 0}$ is weakly continuous. By [EN00, Theorem 5.8], $\operatorname{S}^{*}$ is strongly continuous (we note that more generally this holds on any reflexive Banach space) and accordingly, the map $x\mapsto k_{x}$ is continuous and therefore (locally) integrable. Thus, $\beta$ is well-defined. Since $\Sigma$ is Lipschitz and with bounded support, $\beta$ is Lipschitz. [DZ92, Theorem 7.2] yields existence of the process $(f_{t})_{t\geq 0}$. ∎
Since $\mathcal{M}$ is fully invariant under the HJMM-equation we find that it is invariant under any $\Sigma$ with bounded support fulfilling the regularity conditions outlined in Definition 2.8. In particular, we have $\mathcal{M}\subseteq\operatorname{dom}(\partial_{x})$. This is an adaptation of the classic result
Lemma 4.4.
Let $\mathcal{M}$ be fully invariant under the HJMM-equation. Then we have
$\mathcal{M}\subseteq\operatorname{dom}(\partial_{x}).$
Proof.
Let $h\in\mathcal{M}$ and let $U\subseteq\mathcal{H}$ be an open neighbourhood of $h$. Choose $\Sigma:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ to be of class $C^{1}$ with support in $U$ and $\Sigma(m)\in T_{m}$ for any $m\in U\cap\mathcal{M}$. Since $\mathcal{M}$ is totally invariant we find that $\mathcal{M}$ is invariant for the HJMM-equation with diffusion coefficient $\Sigma$. Hence, [Fil00, Theorem 3] yields that we have
$h\in\mathcal{M}\cap U\subseteq\operatorname{dom}(\partial_{x}).$
Since $h$ was arbitrary we have $\mathcal{M}\subseteq\operatorname{dom}(\partial_{x})$. ∎
In order to prove our main result on the form of invariant manifolds, we consider here the forward rate $f$ and its transformation $h:=\Psi f$. As stated in Remark 2.10, it is natural to construct manifolds as a factor model by defining $f_{t}:=g(\cdot,Y_{t})$ with an appropriate function $g$ and $d$-dimensional diffusion Process $Y$. We will now adapt this method to the transformed process $h$ with an appropriate factor model in the form of a function $g$. The next result shows regularity results of the function $g$. In particular, we find that $g$ is sufficiently smooth for the application of the Itô formula.
Lemma 4.5.
Let $U\subseteq\mathbb{R}^{d}$ be open and connected, $\chi\in C^{2}(U,\mathcal{M})$ and assume that $\partial_{x}\circ\chi\in C(\operatorname{int}(D),\mathcal{H})$ where $D:=\operatorname{dom}(\partial_{x}\circ\chi)$ is its domain. Define
$g(x,y):=\Psi(\chi(y))(x),\quad\text{for }x\geq 0,y\in U.$ (10)
Then $g\in C^{(1,2)}(\mathbb{R}_{+}\times U,\mathbb{R})$.
Proof.
The map $T:a\mapsto 1-\exp(-a)$ is smooth. We have
$g(x,y)=T(\operatorname{J}_{x}(\chi(y))).$
Consequently, $g\in C^{(1,2)}(\mathbb{R}_{+}\times U,\mathbb{R})$ if $l(x,y):=\operatorname{J}_{x}(\chi(y))$ is in $C^{(1,2}(\mathbb{R}_{+}\times U,\mathbb{R})$. Since convergence on $\mathcal{H}$ yields local uniform convergence due to Lemma A.1, we find that it is in $C^{(0,2)}(\mathbb{R}_{+}\times U,\mathbb{R})$. We have $\partial_{x}l(x,y)=\delta_{x}(\chi(y))$ and hence $l\in C^{(1,2)}(\mathbb{R}_{+}\times U,\mathbb{R})$. ∎
We are now prepared to use Itô’s Lemma to characterise the dynamics of the factor model $g$ in view of the dynamics of the HJMM-equation. This results in a drift condition for the invariant manifold induced by $g$ for a model fulfilling the NAFLVR condition for a model specified by a diffusion coefficient.
Lemma 4.6.
Assume the requirements of Lemma 4.5 and let $g$ be the function given therein. Furthermore, let $C\subset U$ be a compact subset and $\sigma\in\mathbb{R}^{d\times d}$ be arbitrary. Let $\varphi\in C^{2}(\mathbb{R}^{d},[0,1])$ be a bump function with $\varphi(y)=1$ for $y\in C$ and $\varphi(y)=0$ for $y\in\mathbb{R}^{d}\setminus U$ and define $\bar{\sigma}(y):=\varphi(y)\sigma$. There exists a measurable function $b^{\sigma}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},y\mapsto b^{\sigma}(y)$ depending on $\sigma$ such that
$\partial_{x}g(x,y)-\partial_{x}g(0,y)+\partial_{x}g(0,y)g(x,y)$ (11)
$=\langle\nabla_{y}g(x,y),b^{\sigma}(y)\rangle+\frac{1}{2}\sum_{i,j=1}^{d}\partial_{y_{i}}\partial_{y_{j}}g(x,y)\bar{\sigma}_{ij}(y)\bar{\sigma}_{ji}(y)$
for any $\sigma\in\mathbb{R}^{d\times d}$, $x\geq 0$ and $y\in U$.
Proof.
Assume the requirements of Lemma 4.5, which yields that $g(x,y):=\Psi(\chi(y))(x)$, $x\geq 0$, $y\in U$ is $C^{(1,2)}(\mathbb{R}_{+}\times U,\mathbb{R})$. Let $\sigma\in\mathbb{R}^{d\times d}$ be arbitrary. Define the function $\tilde{\Sigma}:\mathcal{H}\rightarrow L(\mathbb{R}^{d},\mathcal{H})$ as
[…]
where $\bar{\Sigma}$ is an arbitrary continuous continuation of $D\chi(\chi^{-1}(h))\sigma$. Let now $C\subset U$ be a compact set and $\varphi\in C^{2}(\mathbb{R}^{d},[0,1])$ be a bump function on $C$, as specified in the statement. Set now $\Sigma(h):=\tilde{\Sigma}(h)\varphi(\chi^{-1}(h))$. It is easy to check that, by construction, $\Sigma$ satisfies the requirements of Definition 2.8. Now, since $\mathcal{M}$ is fully invariant under the HJMM-equation, there is a solution $f$ to the HJMM-equation with diffusion coefficient $\Sigma$ such that $\mathbb{P}(f_{t}\in\mathcal{M})=1$ for $t\geq 0$. Define now $Y_{t}:=\chi^{-1}(f_{t})$. By a simple stopping argument, we may restrict ourselves to the case of just a single chart $\chi$. Define $\bar{\sigma}(y):=\varphi(y)\sigma$. By the Inverse Function Theorem, $D\chi^{-1}(h)=(D\chi(\chi^{-1}(h))^{-1}$ and therefore, from Itô’s lemma we get
$dY_{t}=\beta_{t}dt+D\chi^{-1}(f_{t})\Sigma(f_{t})dW_{t}=\beta_{t}dt+\bar{\sigma}(Y_{t})dW_{t},$
where $\beta_{t}:=\text{drift}(f_{t})D\chi^{-1}(f_{t})+\frac{1}{2}\operatorname{Tr}\left(D^{2}\chi^{-1}(f_{t})\left(\Sigma(f_{t}),\Sigma(f_{t})\right)\right)$. Denote the drift of the process $g(x,Y_{t})$ as $\alpha_{t}:=\text{drift}(g(x,Y_{t}))$. By Itô’s lemma, $\alpha$ satisfies
$\alpha_{t}=\langle\nabla_{y}g(x,Y_{t}),\beta_{t}\rangle+\frac{1}{2}\sum_{i,j=1}^{d}\partial_{y_{i}}\partial_{y_{j}}g(x,Y_{t})\bar{\sigma}_{ij}(Y_{t})\bar{\sigma}_{ji}(Y_{t})\quad\text{for all }x,t\geq 0.$
On the other hand, by Lemma 4.2, we find that
$\alpha_{t}=\partial_{x}g(x,Y_{t})-\partial_{x}g(0,Y_{t})g(x,Y_{t})\quad\text{for all }x,t\geq 0.$
Define now for the given $\sigma$ the functions
$\alpha_{t}(y)$ $=\mathbb{E}[\alpha_{t}|Y_{t}=y],$
$b^{\sigma}_{t}(y)$ $=\mathbb{E}[\beta_{t}|Y_{t}=y].$
Then, for all $x\geq 0$ and any $(y,t)\in U\times\mathbb{R}_{+}$ we have $\mathbb{P}^{Y_{t}}$-a.s.
$\partial_{x}g(x,y)-\partial_{x}g(0,y)+\partial_{x}g(0,y)g(x,y)$
$=\langle\nabla_{y}g(x,y),b_{t}^{\sigma}(y)\rangle+\frac{1}{2}\sum_{i,j=1}^{d}\partial_{y_{i}}\partial_{y_{j}}g(x,y)\bar{\sigma}_{ij}(y)\bar{\sigma}_{ji}(y)$
The claim now follows by the same arguments as [KX23, Lemma 2.9]. ∎
Now we can proceed by showing the affine nature of the invariant manifold for the HJMM-equation under our specifications. The next results in essence shows that locally, the image of the factor model $g$ lies in a finite-dimensional affine subspace.
Lemma 4.7.
Let $U\subseteq\mathbb{R}^{d}$ be open and connected and $\chi:U\rightarrow\mathcal{M}$, be a local map, that is $\chi\in C^{2}(U,\mathcal{M})$ with $\operatorname{rank}(D\chi(z))=d$ for any $z\in U$. Define $G:U\rightarrow C(\mathbb{R}_{+},\mathbb{R}),z\mapsto\Psi(\chi(z))$ where $\Psi$ is defined in Equation 3. Then $\operatorname{ran}(G)\subseteq V$ where $V$ is a $d$-dimensional affine subspace of $C(\mathbb{R}_{+},\mathbb{R})$.
The proof of the above Lemma is an adaptation of the proof in [KX23, Proposition 2.14] to our case.
Proof.
Let $C\subset U$ be a compact subset. Without loss of generality we may assume that $0\in C$. Define for $i,j=1,\dots,d$ the symmetric matrices $E_{ij}\in\mathbb{R}^{d\times d}$ as follows
[…]
For a $d\times d$-matrix $a$ and a $C^{2}$ bump function $\varphi$ on $C$ we define, similarly to the proof of Lemma 4.6 the $C^{1}$-function
$\Sigma_{a}:$ $\mathcal{M}\rightarrow L(\mathbb{R}^{d},\mathcal{H}),$
$f\mapsto\varphi(\chi^{-1}(f))D\chi(\chi^{-1}(f))a$
and denote by
$\beta_{a}:$ $\mathcal{M}\rightarrow\mathcal{H},$
$f\mapsto\frac{1}{2}\Sigma_{a}(f)\operatorname{J}_{\cdot}\Sigma_{a}(f)^{*}$
the corresponding HJM-drift coefficient. Since $\mathcal{M}$ is fully invariant under the HJMM-equation we have that it is invariant for solutions to the HJMM-equation with diffusion coefficient $\tilde{\Sigma}_{a}$ where $\tilde{\Sigma}_{a}$ is chosen to be any continuous continuation of $\Sigma_{a}$ outside $\mathcal{M}$. Define $g:\mathbb{R}_{+}\times U\rightarrow\mathbb{R},g:(x,y)\mapsto G(y)(x)$. Lemma 4.5 yields that $g\in C^{(1,2)}(\mathbb{R}_{+}\times U,\mathbb{R})$. By Lemma 4.6, there is a function $b^{a}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, such that
$\partial_{x}g(x,y)-\partial_{x}g(0,y)+\partial_{x}g(0,y)g(x,y)$ (12)
$=\langle\nabla_{y}g(x,y),b^{a}(y)\rangle+\frac{1}{2}\sum_{i,j=1}^{d}\partial_{y_{i}}\partial_{y_{j}}g(x,y)\sigma_{ij}(y)\sigma_{ji}(y)$
Consider now Equation 12 for the two cases $a=\mathbbm{1}_{d}$ and $a=\mathbbm{1}_{d}+E_{ii}$:
$\partial_{x}g(x,y)-\partial_{x}g(0,y)+\partial_{x}g(0,y)g(x,y)$ $=\langle\nabla_{y}g(x,y),b^{\mathbbm{1}_{d}}(y)\rangle+\frac{1}{2}\sum_{j=1}^{d}\partial_{y_{j}}^{2}g(x,y),$
$\partial_{x}g(x,y)-\partial_{x}g(0,y)+\partial_{x}g(0,y)g(x,y)$ $=\langle\nabla_{y}g(x,y),b^{\mathbbm{1}_{d}+E_{ii}}(y)\rangle$
$+\frac{3}{2}\partial_{y_{i}}^{2}g(x,y)+\frac{1}{2}\sum_{j=1}^{d}\partial_{y_{j}}^{2}g(x,y),$
for any $x\geq 0$ and any $y\in U$. Since the left-hand-sides are identical, we then have
$\langle\nabla_{y}g(x,y),b^{\mathbbm{1}_{d}}(y)\rangle+\frac{1}{2}\sum_{j=1}^{d}\partial_{y_{j}}^{2}g(x,y)$
$=\langle\nabla_{y}g(x,y),b^{\mathbbm{1}_{d}+E_{ii}}(y)\rangle+\frac{3}{2}\partial_{y_{i}}^{2}g(x,y)+\frac{1}{2}\sum_{j=1}^{d}\partial_{y_{j}}^{2}g(x,y).$
Hence,
$\partial_{y_{i}}^{2}g(x,y)=\left\langle\nabla_{y}g(x,y),\left(\frac{2}{3}(b^{\mathbbm{1}_{d}}-b^{\mathbbm{1}_{d}+E_{ii}})(y)\right)\right\rangle=\langle\nabla_{y}g(x,y),\eta_{i,i}(y)\rangle,$
where $\eta_{i,i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, $\eta_{i,i}:=\frac{2}{3}(b^{\mathbbm{1}_{d}}-b^{\mathbbm{1}_{d}+E_{ii}})$. $\eta_{i,i}$ is measurable in $y$ and locally bounded by its construction as $b^{\mathbbm{1}_{d}}$ and $b^{\mathbbm{1}_{d}+E_{ii}}$ are locally bounded.
Repeating the same procedure for $\sigma=\mathbbm{1}_{d}$ and $\sigma=\mathbbm{1}_{d}+E_{ij}$ where $i,j=1,\dots,d$ and $i\neq j$, we get analogously to the previous result,
$\langle\nabla_{y}g(x,y),b^{\mathbbm{1}_{d}}(y)\rangle+\frac{1}{2}\sum_{k=1}^{d}\partial_{y_{k}}^{2}g(x,y)$
$=\langle\nabla_{y}g(x,y),b^{\mathbbm{1}_{d}+E_{ij}}(y)\rangle+\partial_{y_{i}}\partial_{y_{j}}g(x,y)+\frac{1}{2}\sum_{k=1}^{d}\partial_{y_{k}}^{2}g(x,y),$
Hence,
$\partial_{y_{i}}\partial_{y_{j}}g(x,y)=\langle\nabla_{y}g(x,y),\left(b^{\mathbbm{1}_{d}}-b^{\mathbbm{1}_{d}+E_{ij}}\right)(y)\rangle=\langle\nabla_{y}g(x,y),\eta_{i,j}(y)\rangle,$
where $\eta_{i,j}:=b^{\mathbbm{1}_{d}}-b^{\mathbbm{1}_{d}+E_{ij}}$, $\eta_{i,j}$ is measurable in $y$ and $\eta_{i,j}$ is locally bounded by its construction as $b^{\mathbbm{1}_{d}}$ and $b^{\mathbbm{1}_{d}+E_{ij}}$ are locally bounded.
Using Equation 11 with $\sigma=\mathbbm{1}_{d}$ and multiplying it by $4$ and subtracting Equation 11 with $\sigma=2\mathbbm{1}_{d}$ yields
$3\left(\partial_{x}g(x,y)-\partial_{x}g(0,y)+\partial_{x}g(0,y)g(x,y)\right)$
$=4\langle\nabla_{y}g(x,y),b^{\mathbbm{1}_{d}}(y)\rangle-\langle\nabla_{y}g(x,y),b^{2\mathbbm{1}_{d}}(y)\rangle$
$=\langle 3\nabla_{y}g(x,y),\gamma(y)\rangle,$
where $\gamma:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},\gamma:=(4b^{\mathbbm{1}_{d}}-b^{2\mathbbm{1}_{d}})/3$. Hence we get for any $i,j=1,\dots,d$, $x\geq 0$, $y\in U$
$\partial_{y_{i}}\partial_{y_{j}}g(x,y)$ $=\langle\nabla_{y}g(x,y),\eta_{i,j}(y)\rangle,$ (13)
$\partial_{x}g(x,y)$ $=\langle\nabla_{y}g(x,y),\gamma(y)\rangle.$ (14)
Due to Equation 13 and Proposition A.6, there exists a twice continuously differentiable function $A:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, where $A(0)=0$ such that
$g(x,y)=g(x,0)+\nabla_{y}g(x,0)A(y),\quad x\geq 0,\ y\in U.$
Define
$V:=g(\cdot,0)+\operatorname{Span}\{\partial_{y_{i}}g(\cdot,0):i=1,\dots,d\}.$
Then $V$ is an affine space of dimension at most $d$. Note that
$G(y)=g(\cdot,y)\in V,\quad y\in U.$
Since $\operatorname{rank}(DG(y))=d$ we find that $I:=\{G(y):y\in U\}$ is a $d$-dimensional $C^{2}$-manifold and the above shows that
$I\subseteq V.$
Consequently, $\dim(V)=d$. ∎
In the next step, we utilise Lemma 4.7 to show that the function $\Psi$ maps a manifold induced by a factor model into an affine subspace.
Proposition 4.8.
There is a $d$-dimensional affine space $V\subseteq C(\mathbb{R}_{+},\mathbb{R})$ such that
$\Psi(\mathcal{M})\subseteq V.$
Proof.
Let $U_{1},U_{2}$ be open and connected and $\chi_{j}:U_{j}\rightarrow\mathcal{M}$ be $C^{2}$-functions with $D\chi_{j}$ full rank everywhere for $j=1,2$ such that $\chi_{1}(U_{1})\cap\chi_{2}(U_{2})\neq\emptyset$, i.e. $\chi_{1},\chi_{2}$ are local maps with a common area. Define $G_{j}:U\rightarrow C(\mathbb{R}_{+},\mathbb{R})$. Note that $G_{1}(U_{1})\cap G_{2}(U_{2})$ is non-empty and a $d$-dimensional manifold. Lemma 4.7 yields affine spaces $V_{1},V_{2}\subseteq C(\mathbb{R}_{+},\mathbb{R})$ of dimension $d$ each such that $G_{j}(U_{j})\subseteq V_{j}$. We find that
$V_{1}\supseteq\left(G_{1}(U_{1})\cap G_{2}(U_{2})\right)\subseteq V_{2}.$
Consequently, $V_{1}=V_{2}$. Since $\mathcal{M}$ is connected we find that $\Psi(\mathcal{M})\subseteq V_{1}$ as claimed. ∎
We are now prepared to provide the proof for our main result.
Proof of Theorem 3.3.
Proposition 4.8 yields a $d$-dimensional affine space $V\subseteq C(\mathbb{R}_{+},\mathbb{R})$ such that
$\Psi(\mathcal{M})\subseteq V.$
Consequently, we have
$\mathcal{M}\subseteq\Psi^{-1}(V).$
Let $h_{0},\dots,h_{d}\in C(\mathbb{R}_{+},\mathbb{R})$ such that
$V=h_{0}+\operatorname{Span}\{h_{1},\dots,h_{d}\}.$
Let $I:=\Psi(\mathcal{M})$. Note that $U$ is connected and open in $V$. Let $U\subseteq\mathbb{R}^{d}$ be open and connected such that the basis representation
$T:\mathbb{R}^{d}\rightarrow V,z\mapsto h_{0}+\sum_{j=1}^{d}z_{j}h_{j}$
satisfies $T(U)=I$. We find that
$\mathcal{M}=\Psi^{-1}(I)=\Psi^{-1}(T(U)).$
The claim follows. ∎
Proof of Theorem 3.6.
Let $\mathcal{H}^{0}$ and $\mathcal{M}\subseteq\mathcal{H}^{0}$ be defined as in the statement.
  • i) $\Rightarrow$ ii):
    Theorem 3.3 yields the asserted manifold $\mathcal{M}$ for some $C^{2}$ functions $c:\mathbb{R}_{+}\rightarrow\mathbb{R}$ and $u:\mathbb{R}_{+}\rightarrow\mathbb{R}$. Using the results of [Fil23, Proposition 3.2] we obtain a matrix $M\in\mathbb{R}^{(d+1)\times(d+1)}$ such that $c$ and $u$ assume the asserted form.
  • ii) $\Rightarrow$ i):
    Let $\mathcal{B}$ and $\Lambda_{\text{HJMM}}$ be as in Definition 2.8 and $f_{0}\in M$, $(\beta,\Sigma)\in\Lambda_{\text{HJMM}}$. Lemma 4.3 yields that there is an $\mathcal{H}$-valued solution $f$ to the HJMM-equation $(\beta,\Sigma)$ with starting value $f_{0}$. We need to show that $f$ is $\mathcal{M}$-valued.
    Consider the map $\Psi$ from Equation 3 and define $V=\Psi(\mathcal{M})$, as well as $h_{t}:=\Psi f_{t}$ for all $t\geq 0$. Define
    $\mathcal{B}^{h}=\{\Sigma^{h}:\mapsto(1-h)\operatorname{J}_{\cdot}\Sigma(\Psi^{-1}h)\;|\;\Sigma\in\mathcal{B}\}$
    and consider the parameter set
    $\Lambda^{h}:=\{(\beta^{h},\Sigma^{h})\;|\;\beta^{h}(h):=(h-1)\delta_{0}(\Psi^{-1}h),\Sigma^{h}\in\mathcal{B}^{h}\}.$
    Due to Lemma 4.2, if $f$ is a solution to the HJMM-equation with $(\beta,\Sigma)\in\Lambda_{\text{HJMM}}$, then $h$ is a solution to the SPDE
    $dh_{t}=(\partial_{x}h_{t}+\beta^{h}(h_{t}))dt+\Sigma^{h}(h_{t})dW_{t}$ (15)
    with $(\beta^{h},\Sigma^{h})\in\Lambda^{h}$. Indeed, since $\Psi$ is injective, in order to show that $\mathcal{M}$ is fully invariant under the HJMM-equation, it is sufficient to show that $V$ is fully invariant under Equation 15 for $\Lambda^{h}$. We may therefore verify the conditions of Remark 2.9 for the manifold $V$ to prove the assertion. Using the results of Proposition 4.8, we see that
    $V=\bigg\{c+\sum_{i=1}^{d}u_{i}z_{i}\;\bigg|\;z\in U\bigg\}.$
    Since $c$ and $u$ are real analytic functions, this easily implies $V\subseteq\text{dom}(\partial_{x})$, that is 2.9.ii)a. Now, consider the pair $(\beta^{h},\Sigma^{h})\in\Lambda^{h}$. Since $\operatorname{J}_{\cdot}$ is a linear operator, $\Sigma^{h}(h)\in L(\mathbb{R}^{d},T_{h})$ for any $h\in V$, which is directly 2.9.ii)b of. To verify the third condition, we observe that $V$ is an affine manifold. Due to [Fil00, Theorem 4], this reduces Condition 2.9.ii)c to $\partial_{x}h+\beta^{h}(h)\in T_{h}$ for any $h\in T_{h}$. To show that this is true, define $r:=\delta_{0}(\Psi^{-1}h)$. Due to Lemma 4.1, we find that $r=h^{\prime}(0)$. Using the particular form of the functions $c$ and $u$, we therefore have $r=\langle-Me_{1},(1,z)\rangle$ for $z\in U$. We may now compute
    $\partial_{x}h+\beta^{h}(h)=\partial_{x}h+(h-1)r$
    $=\partial_{x}\langle(\mathbbm{1}_{d+1}-e^{\cdot M})e_{1},(1,z)\rangle+r\left(\langle(\mathbbm{1}_{d+1}-e^{\cdot M})e_{1},(1,z)\rangle-1\right)$
    $=-\langle Me^{\cdot M}e_{1},(1,z)\rangle-r\langle e^{\cdot M}e_{1},(1,z)\rangle$
    $=-\langle(M-r\mathbbm{1}_{d+1})e^{\cdot M}e_{1},(1,z)\rangle=-\langle e^{\cdot M}e_{1},(M^{\top}+\mathbbm{1}_{d+1})(1,z)\rangle$
    $=-\langle e^{\cdot M}e_{1},M^{\top}(1,z)-\langle e_{1},M^{\top}(1,z)\rangle(1,z)\rangle$
    Observe now that the vector $v:=M^{\top}(1,z)-\langle e_{1},M^{\top}(1,z)\rangle(q,z)\rangle$ satisfies $v_{1}=M_{1,1}+\sum_{i=1}^{d}M_{i+1,1}z_{i}-(M_{1,1}-\sum_{i=1}^{d}M_{i+1,1}z_{i})=0$. This implies $\partial_{x}h+\beta^{h}(h)\in\text{span}\{u_{1},\dots,u_{d}\}$. Since $V$ is an affine manifold, its tangent space satisfies $T_{h}=V_{0}$, where $V_{0}=\text{span}\{u_{1},\dots,u_{d}\}$, the vector space obtained by shifting the affine space to the origin. This implies that Condition 2.9.ii)c is satisfied and therefore proves the assertion.
Appendix A Technical tools
Lemma A.1.
We have $\mathcal{H}\subseteq C(\mathbb{R}_{+},\mathbb{R})$. Moreover, convergence in $\mathcal{H}$ implies local uniform convergence.
Proof.
Observe that $f(x)=\delta_{0}\operatorname{S}_{x}f$ for $x\geq 0$. Since $\operatorname{S}$ is strongly continuous we find that $x\mapsto\operatorname{S}_{x}f$ is continuous and, hence, $\mathcal{H}\subseteq C(\mathbb{R}_{+},\mathbb{R})$. Moreover, let $(f_{n})_{n\in\mathbb{N}}$ be a convergent sequence in $\mathcal{H}$ with limit $f_{\infty}\in\mathcal{H}$. We have for $x_{1}\geq 0$
$\sup_{x\in[0,x_{1}]}|f_{n}(x)-f_{\infty}(x)|\leq\sup_{x\in[0,x_{1}]}\|\operatorname{S}_{x}\|\|\delta_{0}\||f_{n}-f_{\infty}|\rightarrow 0$
as $n\rightarrow\infty$ because [EK09, Ch. 1, Proposition 1.2] yields that $\sup_{x\in[0,x_{1}]}\|\operatorname{S}_{x}\|<\infty$. ∎
Lemma A.2.
Let $x\geq 0$. The linear functional $\operatorname{J}_{x}:\mathcal{H}\rightarrow\mathbb{R}$ given by
$\operatorname{J}_{x}(f):=\int_{0}^{x}f(s)ds$
is defined for all $f\in\mathcal{H}$ and is continuous and linear. Moreover, the map $x\mapsto\operatorname{J}_{x}(f)$ is in $C^{1}(\mathbb{R}_{+},\mathbb{R})$ and
$\operatorname{J}:\mathbb{R}_{+}\times\mathcal{H}\rightarrow\mathbb{R},\quad(x,f)\mapsto\operatorname{J}_{x}(f)$
is continuous.
Proof.
Lemma A.1 yields that $f\in\mathcal{H}$ is continuous. Hence, $\operatorname{J}_{x}(f)$ is well defined and $x\mapsto\operatorname{J}_{x}(f)$ is continuously differentiable. Furthermore, if $f_{n}\rightarrow f_{\infty}$ in $\mathcal{H}$ and $x_{n}\rightarrow x_{\infty}$ in $\mathbb{R}_{+}$, then $f_{n}\rightarrow f_{\infty}$ locally uniformly on $0,\sup_{n\in\mathbb{N}}x_{n}]$ and, hence we find that
$\operatorname{J}(x_{n},f_{n})\rightarrow\operatorname{J}(x_{\infty},f_{\infty}).$
In particular, $\operatorname{J}_{x}$ is continuous and linear. ∎
We now provide the classical result that the generator $\partial_{x}$ of $\operatorname{S}$ is the derivative operator. A formal proof is given for completeness.
Lemma A.3.
We have
$\operatorname{dom}(\partial_{x})=\{f\in\mathcal{H}\cap C^{1}(\mathbb{R}_{+},\mathbb{R}):f^{\prime}\in\mathcal{H}\}$
and $\partial_{x}f=f^{\prime}$ for $f\in\operatorname{dom}(\partial_{x})$, where $f^{\prime}$ denotes the derivative of the function $f$.
Proof.
Let $f\in\operatorname{dom}(\partial_{x})$. Then $\partial_{x}f=\lim_{h\searrow 0}\frac{\operatorname{S}_{h}f-f}{h}$. We find that
$(\partial_{x}f)(x)=\delta_{x}(\partial_{x}f)=\lim_{h\searrow 0}\frac{f(x+h)-f(x)}{h}=f^{\prime}(x+)$
where $f^{\prime}(\cdot+)$ denotes the right-derivative of $f$. Since $f^{\prime}(\cdot+)\in\mathcal{H}\subseteq C(\mathbb{R}_{+},\mathbb{R})$ we find for $x>h>0$ that
$\frac{f(x-h)-f(x)}{-h}=\delta_{x-h}\left(\frac{\operatorname{S}_{h}f-f}{h}\right)\rightarrow\delta_{x}\left(\partial_{x}f\right)=f^{\prime}(x+)$
as $h\searrow 0$ because $\frac{\operatorname{S}_{h}f-f}{h}\rightarrow f^{\prime}(\cdot+)$ locally uniformly by Lemma A.1. Consequently, the left-derivative equals the right-derivative and $f^{\prime}(\cdot+)=f^{\prime}\in\mathcal{H}\subseteq C(\mathbb{R}_{+},\mathbb{R})$. Hence,
[…]
Now, let $f\in\mathcal{H}\cap C^{1}(\mathbb{R}_{+},\mathbb{R})$ with $f^{\prime}\in\mathcal{H}$ and we show that $f\in\operatorname{dom}(\partial_{x})$. We have $h\mapsto\operatorname{S}_{h}f^{\prime}$ is continuous, hence, locally integrable. Thus, $\int_{0}^{h}\operatorname{S}_{u}f^{\prime}du<\infty$ and we have
$\delta_{x}\left(\int_{0}^{h}\operatorname{S}_{u}f^{\prime}du\right)=\int_{0}^{h}f^{\prime}(u+x)du=f(x+h)-f(x)=\delta_{x}(\operatorname{S}_{h}f-f),$
which shows that $\int_{0}^{h}\operatorname{S}_{u}f^{\prime}du=\operatorname{S}_{h}f-f$. The fundamental theorem of calculus yields
$f^{\prime}=\lim_{h\searrow 0}\frac{1}{h}\int_{0}^{h}\operatorname{S}_{u}f^{\prime}du=\lim_{h\searrow 0}\frac{\operatorname{S}_{h}f-f}{h},$
which shows that $f\in\operatorname{dom}(\partial_{x})$ and $\partial_{x}f=f^{\prime}$. ∎
Lemma A.4.
Let $V\subseteq\mathfrak{F}(\mathbb{R}_{+},\mathbb{R})$ be a finite-dimensional vector space of functions. Let $g\in V$ be such that $g^{n}\in V$ for any $n\in\mathbb{N}$. Then $g$ is constant
Proof.
Since $V$ is finite-dimensional, there is a minimal $N\in\mathbb{N}$, so that $g^{1},\dots,g^{N}$ are linearly dependent. Therefore, there exist $c_{1},\dots,c_{N-1}\in\mathbb{R}$, so that
$g^{N}=\sum_{i=1}^{N-1}c_{i}g^{i}.$
Let now $\mathcal{U}:=\text{span}\{g^{1},\dots g^{N-1}\}$. Then $\mathcal{U}$ is an algebra of functions. Define the linear operator $M_{g}:\mathcal{U}\rightarrow\mathcal{U}$, $h\mapsto gh$. Let $h$ be an eigenvector of $M_{g}$. We find $M_{g}h=\lambda h$ for some $\lambda\in\mathbb{C}$. This implies
$0=(M_{g}-\lambda)h=(g-\lambda)h.$
Thus, $\{h\neq 0\}\subseteq\{g=\lambda\}$. Since this is true for any eigenvector $h$, we have $\mathrm{ran}(g)\subseteq\text{EV}(M)$, where $\text{EV}(M)$ denotes the set of eigenvalues of the operator $M$. ∎
Lemma A.5.
Let $\mathcal{M}$ be linear-rational in the sense of Definition 3.1 and $h_{0},h_{1}\in\mathcal{H}$. Assume that $f:\mathbb{R}_{+}\rightarrow\mathcal{H}$, $t\mapsto h_{0}+th_{1}$ is $\mathcal{M}$-valued. Then $h_{1}\equiv 0$.
Proof.
Since $f$ is continuous and with values in $\mathcal{M}$, we find $\gamma:[0,1]\rightarrow\mathbb{R}^{d}$ such that
$f(t)(x)=\frac{c^{\prime}(x)+\langle u^{\prime}(x),\gamma(t)\rangle}{1-\left(c(x)+\langle u(x),\gamma(t)\rangle\right)}\qquad\text{for all }t\in[0,1],x\geq 0.$
Define now $H_{0}(x):=\operatorname{J}_{x}h_{0}$, $H_{1}(x):=\operatorname{J}_{x}h_{1}$ and we find
$\exp\left(-H_{0}(x)-tH_{1}(x)\right)=\exp\left(-\int_{0}^{x}f(t)(s)ds\right)=1-\left(c(x)+\langle u(x),\gamma(t)\rangle\right)\in V,$
where $V$ is the vector space spanned by $\{1-c,u_{1},\dots,u_{d}\}$.
Define $g(x):=\exp\left(-H_{1}(x)\right)$ and the finite dimensional vector space $\mathcal{U}:=\{\exp\left(H_{0}\right)v:v\in V\}$. We observe that $g^{n}=\exp\left(-nH_{1}\right)\in\mathcal{U}$ by construction. Lemma A.4 yields that $g$ is constant. Consequently, $H_{1}$ is constant and therefore $h_{1}\equiv 0$ as claimed. ∎
Proposition A.6.
Let $U\subseteq\mathbb{R}^{d}$ be open. Let $\eta_{i,j}:U\rightarrow\mathbb{R}^{d}$ be measurable and locally bounded with $\eta_{i,j}=\eta_{j,i}$ for any $i,j=1,\dots,d$.
Then there is a twice continuously differentiable function $A:U\rightarrow\mathbb{R}^{d}$ such that for any twice continuously differentiable function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ with
$\partial_{i}\partial_{j}g(y)=\langle\nabla g(y),\eta_{i,j}(y)\rangle,\quad i,j=1,\dots,d,y\in U,$
one has
$g(y)=g(0)+\langle\nabla g(0),A(y)\rangle,\quad y\in\mathbb{R}^{d}.$
Proof.
Without loss of generality we may assume that $0\in U$.
Define
$\mathcal{S}$ $=\{f\in C^{2}(U,\mathbb{R}):\partial_{i}\partial_{j}f(y)=\langle\nabla f(y),\eta_{i,j}(y)\rangle,\text{ for any }i,j=1,\dots,d,y\in U\},$
$\mathcal{S}_{0}$ $=\{f\in\mathcal{S}:f(0)=0\}.$
Note that $\mathcal{S}$, $\mathcal{S}_{0}$ are vector spaces and for $f\in\mathcal{S}$ one has $h:=f-f(0)\in\mathcal{S}_{0}$ and
$f=f(0)+h,$
i.e. $\mathcal{S}$ is the direct sum $\mathcal{F}_{0}\oplus\mathcal{S}_{0}$ where $\mathcal{F}_{0}$ is the space of constant functions.
Let $h\in\mathcal{S}_{0}$ with $\nabla h(0)=0$. We show that $h=0$. To this end let $R>0$. By assumption on $\eta$ there is a constant $C\geq 0$ such that $\eta_{i,j}$ is bounded on the ball with radius $R$ by $C$ for any $i,j=1,\dots,d$. We find for $x\in\mathbb{R}^{d}\setminus\{0\}$ with $|x|\leq R$ that
$|\nabla h(x)|=\left|\int_{0}^{1}D(\nabla h)(tx)xdt\right|\leq Cd^{2}\int_{0}^{1}\left|\nabla h(tx)\right||x|dt.$
Thus, Grönwall’s lemma yields that
$|\nabla h(x)|\leq|\nabla h(0)|\exp(|x|Cd^{2})=0,$
for any $x\in\mathbb{R}^{d}$ with $|x|\leq R$. Consequently, $\nabla h=0$ which yields that $h$ is constant. Since $h(0)=0$ we find that $h=0$.
Define $\Theta:\mathcal{S}_{0}\rightarrow\mathbb{R}^{d},f\mapsto\nabla f(0)$. By the above we have that $\Theta$ is an injective linear map. Consequently, $l:=\dim(\mathcal{S}_{0})\leq d$. Let $f_{1},\dots,f_{l}$ be maximal linear independent in $\mathcal{S}_{0}$ such that $b^{j}:=\Theta(f_{j})$, $j=1,\dots,l$ is an orthonormal system with respect to the standard scalar product. There is an orthogonal transformation $T:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ with $Tb^{j}=e_{j}$ for any $j=1,\dots,l$. We define
$f(y)$= $(f_{1}(y),\dots,f_{l}(y),0,\dots,0),$
$A(y)$= $T^{\top}(f(y)),$
for any $y\in\mathbb{R}^{d}$. Note that $A:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is twice continuously differentiable and $A(0)=0$.
Moreover, we have for $j=1,\dots,l$ and $y\in\mathbb{R}^{d}$ that
$f_{j}(0)+\langle\nabla f_{j}(0),A(y)\rangle=\langle b^{j},T^{\top}(f(y))\rangle=\langle T(b^{j}),f(y)\rangle=\langle e_{j},f(y)\rangle=f_{j}(y).$
By linearity we find that
$h(0)+\langle\nabla h(0),A(y)\rangle=h(y),$
for any $y\in U$, $h\in\mathcal{S}_{0}$. The claim follows. ∎